OpenStudy (moongazer):

3csc^2(3pi/2)+2sin(3pi/2) - (/3sin3pi/)/2 is the answer equal to 1?

OpenStudy (anonymous):

To clarify, is it: \[3cosec^2\left(\frac{3\pi}{2}\right)+2sin\left(\frac{3\pi}{2}\right)-\frac{1}{3}sin\left(\frac{3\pi}{2}\right)?\]

OpenStudy (moongazer):

3cosec2(3π2)+2sin(3π2) part is correct but the last should be -(/3sin^3pi/)/2 / on 3sin^3pi means absolute value

OpenStudy (anonymous):

\[\frac{|3sin^3(\pi)|}{2}\]?

OpenStudy (moongazer):